# Distributed Stochastic Gradient Descent with Staleness: A Stochastic Delay Differential Equation Based Framework

### v1

@@ III-C The Convergence Speed Given Gradient Staleness
Remark:  Assuming an $L$-Lipschitz continuous objective function $f(\boldsymbol{x})$, the step size must be chosen as $\eta<\frac{\pi}{2L\breve{\tau}}$ to avoid unbounded oscillations. A more conservative choice $\eta<\frac{1}{eL\breve{\tau}}$ ensures that no oscillation occurs. In this case, a small degree of gradient staleness accelerates the SGD algorithm, which aligns with the conclusion in [30], and [31] in which quantization error can be exploited to accelerate the model convergence. Moreover, an appropriate staleness accelerates SGD the convergence in the direction of eigenvectors with small eigenvalues and hinders others with large eigenvalues. The above discussion aligns with the observation made by [14] that a stale gradient may contribute to a faster convergence rate but a higher error floor. In some existing literature [19, 18], the stale-aware step size is inversely proportional to stale, which also aligns with our analysis. However, step size must be carefully chosen with further consideration of the 2-norm of the Hessian matrix of the objective function, to avoid unbounded oscillations.
Next, we proceed with the following discussion on the approximation of the SDDE with a small constant staleness. Recall that we have $\mathrm{d}x(t)=-vx(t-\tau)\mathrm{d}t+\sigma\mathrm{d}B(t)$. Further, let $t_{2}=t_{1}-\tau$, and the random variable $X_{t_{2}}=x(t_{2})$ denote the state value that $x(t)$ takes at time $t_{2}$. In this way, $f^{\prime}(x(t_{2}))=vX_{t_{2}}$. Additionally, $X_{t_{2}}$ is Gaussian distributed given $x(t_{1})=X_{t_{1}}$ by the definition of an OU process. In this way, it holds that $\mathbb{E}\{X_{t_{2}}|x(t_{1})=X_{t_{1}}\}=X_{t_{1}}e^{v\tau},$ and $\text{Var}\{X_{t_{2}}|x(t_{1})=X_{t_{1}}\}=\frac{\sigma^{2}}{2}(e^{2v\tau}-1).$ Also, it holds that $\text{Var}\{X_{t_{2}}|x(t_{1})=X_{t_{1}}\}\approx v\tau\sigma^{2}\text{ for }\tau\ll 1$. By this means, we conclude that $X_{t_{2}}\sim\mathcal{N}\left(X_{t_{1}}e^{v\tau},\frac{\sigma^{2}}{2}(e^{2v\tau}-1)\right).$ Therefore, we have $\mathrm{d}x(t)=-ve^{v\tau}x(t)\mathrm{d}t+\sqrt{\frac{e^{2v\tau}+1}{2}}\sigma\mathrm{d}B(t).$ When $\tau\ll 1$, it holds that $\mathrm{d}x(t)=-v(1+v\tau)x(t)\mathrm{d}t+{(1+v\tau/2)}\sigma\mathrm{d}B(t).$
@@
In the following, we present the performance analysis when the gradient staleness is a uniform random variable. In particular, we focus on the uniformly distributed gradient staleness, which can arise in event-triggered SGD. More specifically, consider the delay differential equation given below:
@@
For $\zeta\in(0,\frac{\varpi^{2}}{e^{-\varpi}-1})$, there are two negative real solutions, in which it holds that $|\lambda_{0}|<|\lambda_{1}|$. Note that $\lambda_{0}$ is the dominant root. Further, we have the following approximation for the dominant root $\lambda_{0}$ in the following corollary.
Corollary 5
It holds that $\lambda_{0}\approx\frac{2}{\zeta-2}$ for $\zeta\ll 1$.
Proof:
By applying the second order Taylor expansion, it follows that $\zeta\lambda^{2}\approx-\zeta\lambda+\frac{\zeta^{2}\lambda^{2}}{2}$. By this means, we arrive at Corollary 5. ∎
@@ IV Parameter Optimization for Distributed SGD
IV-A Parameter Optimization for Asynchronous SGD
In this subsection, we present the probability distribution of the update interval, also referred to as the run time[14], and step staleness in $B$-ASGD to bridge the asynchronous SGD algorithm and the SDDE-based continuous approximation. Recall that in $B$-ASGD, the central server collects the first $B$ gradients while the other clients do not halt the computation of their gradient. For the sake of discussion, we assume infinite communication bandwidth in this subsection, in which the communication delays are negligible compared to the computational delay. In this case, the gradient delay or the continuous-time gradient staleness equals the computational delay. The computational delay of each client is assumed to follow a general distribution, in which they are assumed to have a common expectation, i.e., $\mathbb{E}\{T_{\kappa}\}=1/\nu$. Let $N_{p}(t)$ be the aggregated arrival process of gradients to the central server in $B$-ASGD. To ensure the validity of the Poisson approximation, it is assumed that $B$ is much smaller than $K$ in this section. Let $Y\sim\text{Exp}(\nu)$ denote a random variable following an exponential distribution with a mean $1/\nu$. By [43, 44, 45], we have the following lemma.
@@ IV-A Parameter Optimization for Asynchronous SGD
Given that the central server has collected $i\in\{0,1,...,B-1\}$ gradients since the last update of the global model, $N_{p}(t)$ can be approximated as a Poisson process with rate $(N-i)\nu$.
In this way, we have the following theorem, which does not rely on exponentially distributed computation time.
Theorem 2
The update intervals in $B$-ASGD is given by
@@
in which $I^{(i)}\sim\text{Exp}((K-i+1)\nu)$ and $\mathbb{E}\{I\}=\frac{1}{v}\sum_{i=K-B+1}^{K}\frac{1}{i}$. Also, it holds that
@@
By the Poisson approximation of the aggregated arrival process, the update interval between the arrival of the $(i-1)$th and $i$th gradients is exponentially distributed with rate $(K-i+1)\nu$ for $i\in\{1,...,B\}$. By this means, we get Eq. (23). Next, we have Eq. (24) by the Euler–Maclaurin formula. ∎
@@
The expectation of the step staleness in $B$-ASGD without gradient dropout is given by [^12] When $B=1$, this theorem reduces to the pure asynchronous SGD case, which was presented by Eq. (18) in [15].
@@
It is observed that in $B$-ASGD, the reduction of step staleness is achieved by decreasing the update frequency. Next, we present a brief discussion on the difference between $B$-ASGD and $B$-batch ASGD. Similar to Theorem 2, the update interval follows an Erlang distribution, denoted by $\text{Erlang}(B,\nu)$, with mean $B/\nu$ in $B$-batch ASGD. Further, following the discussion in Theorem 3, it holds that $\mathbb{E}\{\breve{\tau}\}=K-1$ in $B$-batch ASGD.
To further reduce the step staleness, the client could choose to stop the computation of the current gradient according to a continuous-time or discrete-time staleness threshold. Next, we first present the following discussion on $B$-ASGD with a continuous-time staleness threshold. In particular, we assume each client’s computational delay follows an i.i.d. distribution, denoted by $T$. Let $F_{T}(\cdot)$ be the cumulative distribution function of $T$. In this context, let $p=F_{T}(\hat{\tau}_{th})$ be the probability that a computation of gradient is not discarded due to timeout. Further, $\bar{p}=1-p$. Let $\tilde{T}$ be the truncated computation time, in which $F_{\tilde{T}}(s)=\frac{F_{T}(s)}{p}$ if $s<\hat{\tau}_{th}$ and $F_{\tilde{T}}(s)=0$ for $s>\hat{\tau}_{th}$. In this case, $\mathbb{E}\{\tilde{T}\}=\frac{1}{p}\int_{s=0}^{\hat{\tau}_{th}}s{\rm d}{F_{T}(s)}$. Next, we have the following proposition on the inter-update time of the $B$-ASGD with a continuous-time staleness threshold.
@@
Next, we have the following theorem for the step staleness in $B$-SGD with a continuous-time staleness threshold.
Corollary 8
The expectation of the step staleness in $B$-ASGD with a continuous time staleness threshold $\hat{\tau}_{th}$ is approximately given by $\mathbb{E}\{\breve{\tau}\}\approx(K-B)\left(1-\frac{\bar{p}\hat{\tau}_{th}}{\bar{p}\hat{\tau}_{th}+\int_{s=0}^{\hat{\tau}_{th}}s{\rm d}{F_{T}(s)}-\beta p\mathbb{E}\{\tilde{I}\}}\right)$. [^13] Here, $\beta\in(0,1)$ is a constant introduced to offset the non-Poissonity of the global model update process.
Next, we turn our attention to $B$-ASGD with a discrete-time gradient staleness threshold $\breve{\tau}_{th}$. In particular, if the global model has been updated $\breve{\tau}_{th}$ times before a client finishes its computation of the current gradient, the client shall stop the computation of the current gradient and starts the computation of a new one. In this case, the step staleness of each gradient is no greater than $B\breve{\tau}_{th}$. With memoryless computation time, we have the following proposition.
@@
$\text{Pr}\{\breve{\tau}=mB\}=\frac{q\bar{q}^{m}}{1-\bar{q}^{\breve{\tau}_{th}+1}}\text{ for }\breve{\tau}=0,B,...,B\breve{\tau}_{th}$ (27)
@@
IV-B Parameter Optimization for Pure Asynchronous SGD
In this subsection, we focus on pure asynchronous SGD, or $1$-ASGD, for deeper insights, in which $B=1$ and $\hat{\tau}_{th}=\infty$. Here we relax the assumption of identically distributed computation time across each client. Let $T_{\kappa}$ be the continuous-time random variable corresponding to the computational delay of the $\kappa$th worker. Further, it is assumed that $0<\mathbb{E}\{T_{\kappa}\}<\infty$. Let $\nu_{\kappa}=\frac{1}{\mathbb{E}\{T_{\kappa}\}}$ and $\nu_{p}=\sum_{\kappa=1}^{K}\nu_{\kappa}$. The sequence of the computational delay from Worker $\kappa$ is denoted by $\left(T_{\kappa}^{(i)}\right)_{i\geq 1}$. The update time of each worker can therefore be denoted by $J^{(n)}_{\kappa}=\sum_{i=1}^{n}T_{\kappa}^{(i)}$. Let us define the following stochastic process $N_{p}^{\kappa}(t)=\sum_{n=1}^{\infty}\mathds{1}\left\{J^{(n)}_{\kappa}\leq t\right\}$ representing the number of updates from Worker $\kappa$ by time $t$. Furthermore, the rate of $N_{p}^{\kappa}(t)$ is $\nu_{\kappa}$. In this context, the number of arrived gradients at the parameter server is given by the superposition of $N_{p}^{\kappa}(t)$ for $\kappa=1,2,...,K$. More specifically, we have
$\vspace{-1.5ex}N_{p}(t)=\sum_{\kappa=1}^{K}N_{p}^{\kappa}(t).$ (29)
From Theorem 2, we have the following corollary.

### v2

@@ III-C The Convergence Speed Given Gradient Staleness
Remark:  Assuming an $L$-Lipschitz continuous objective function $f(\boldsymbol{x})$, the step size must be chosen as $\eta<\frac{\pi}{2L\breve{\tau}}$ to avoid unbounded oscillations. A more conservative choice $\eta<\frac{1}{eL\breve{\tau}}$ ensures that no oscillation occurs. In this case, a small degree of gradient staleness accelerates the SGD algorithm, which aligns with the conclusion in [32], and [63] in which quantization error can be exploited to accelerate the model convergence. Moreover, an appropriate staleness accelerates the convergence of SGD in the direction of eigenvectors with small eigenvalues and hinders others with large eigenvalues. The above discussion aligns with the observation made by [16] that a stale gradient may contribute to a faster convergence rate but a higher error floor. In some existing literature [20, 19], the stale-aware step size is inversely proportional to stale, which also aligns with our analysis. However, step size must be carefully chosen with further consideration of the 2-norm of the Hessian matrix of the objective function, to avoid unbounded oscillations.
In the above discussion, it has been revealed that the convergence rate of the SGD is determined by $v$ and gradient staleness in SDDE $\tau$. In the following, we shall present the stationary distribution of ASGD with gradient staleness by [64].
Proposition 3
When $0<v\tau<\frac{\pi}{2}$, the stationary distribution of the SDDE $\mathrm{d}x(t)=-{v}x(t-\tau)\mathrm{d}t+\sigma\mathrm{d}B(t)$ is Gaussian with zero mean and variance
$\sigma_{\pi}=\frac{\sigma^{2}}{2v}\left(\frac{1+\sin(v\tau)}{\cos(v\tau)}\right).$ (31)
for $0<v\tau<\frac{\pi}{2}$.
Using a Taylor expansion for $\tau\ll 1$, we obtain $\sigma_{\pi}=\frac{\sigma^{2}}{2v}\left(1+v\tau\right)+\mathcal{O}(\tau^{2})$. Here, the result Eq. (31) aligns with our earlier analysis in Eq. (27), where ASGD diverges when $v\tau>\frac{\pi}{2}$. Additionally, ASGD is found to have a larger variance in its stationary distribution compared to synchronous SGD without gradient staleness. This implies that ASGD has a greater steady-state error, defined as the difference between the desired value and the actual value of a system once the response has reached the steady state. To improve stability and convergence in practical applications, one could reduce the level of asynchrony or increase the mini-batch size, particularly when the algorithm is approaching a local minimum.
In the following discussion, we are interested in the characteristic roots of the SDDE when $v\leq 0$. This analysis is particularly relevant for understanding the behavior of ASGD near saddle points. When $v<0$, there are two characteristic roots, $\tilde{\lambda}_{0}=\frac{W_{0}(-v\tau)}{\tau}$ and $\tilde{\lambda}_{1}=\frac{W_{1}(-v\tau)}{\tau}$. In this case, $\tilde{\lambda}_{0}$ is real and positive, while $\text{Re}\left\{\tilde{\lambda}_{0}\right\}<0$. In the following, we primarily focus on the dominant characteristic root $\tilde{\lambda}_{0}$ which has a positive real part. For brevity, we denote $\tilde{\lambda}_{0}$ by $\tilde{\lambda}$.
Proposition 4
If $v<0$, the real and positive characteristic root of Eq. (26) is given by
$\tilde{\lambda}(\tau)=\frac{W_{0}(-v\tau)}{\tau}.$ (32)
Recall that $W_{0}(\cdot)$ is the Lambert W function. Note that to differentiate from the case when $v>0$, we denote the characteristic root as $\tilde{\lambda}$ specifically when $v<0$. Next, we have the following corollary.
Corollary 3
It holds that
$\lim_{\tau\to 0}\tilde{\lambda}(\tau)=-v.$ (33)
Proof:
Since $\frac{\rm d}{\rm dt}W_{0}(0)=1$, we obtain Eq. (33) by L’Hôpital’s rule. ∎
Further, we have the following corollary revealing the monotonicity of $\tilde{\lambda}(\tau)$.
Corollary 4
When $v<0$, the function $\tilde{\lambda}=\frac{W_{0}(-v\tau)}{\tau}$ is a decreasing function of $\tau$ on the interval $[0,\infty)$.
Proof:
The derivative of $\tilde{\lambda}$ with respect to $\tau$ is given by
$\frac{\rm d}{\rm d\tau}\tilde{\lambda}(\tau)=\frac{W_{0}(-v\tau)\left(-\frac{v}{W_{0}(-v\tau)+1}-1\right)}{\tau^{2}}.$
For $v<0$, $W_{0}(-v\tau)>0$ since $-v\tau>0$. The term $-\frac{v}{W_{0}(-v\tau)+1}-1$ is negative, leading to $\tilde{\lambda}^{\prime}(\tau)<0$. Thus, $\tilde{\lambda}$ is decreasing with respect to $\tau$. ∎
Therefore, we have the following proposition, which demonstrates that gradient staleness slows down the escape from a saddle point.
Proposition 5
It holds that $\tilde{\lambda}<-v$ when $v<0$.
Next, we have the following approximations on the characteristic root.
Proposition 6
If $\tau\ll 1$, the characteristic root can be approximated as $\tilde{\lambda}=-v(1+v\tau)+\mathcal{O}(\tau^{2})$.
Proof:
Since $W_{0}(s)=s-\frac{s^{2}}{2}+\mathcal{O}(s^{3})$ for $s\ll 1$ by Taylor expansion, we arrive at the conclusion. ∎
As stated in Assumption 1, the objective function is assumed to be twice differentiable, which motivates further insightful discussions highlighting the impact of the gradient staleness on the behavior of the SGD algorithm based on the positive definiteness of the Hessian matrix. As presented above, a small degree of gradient staleness accelerates convergence, but a large degree of gradient staleness causes divergence near critical points where the eigenvalues of the Hessian matrices are positive. For the sake of completeness, we next highlight the impact of gradient staleness on the behavior of SGD near critical points, where some eigenvalues of the Hessian matrix can be negative, to provide further insights. In particular, the presence of gradient staleness results in a slower escape from the critical point with only negative eigenvalues. This occurs because the effective gradient magnitude is scaled by a factor of $(1+v\tau)$ when $\tau\ll 1$ and $v<0$. Near a saddle point, the escape speed depends on both the magnitude of gradient staleness and the Hessian matrix of the objective function. Gradient staleness can either accelerate or decelerate the escape from the saddle points. Consider a two-dimensional objective function, where the eigenvalues of the Hessian matrix are $v_{1}>0$ and $v_{2}<0$ and $\tau\ll 1$. When $|v_{2}|<|v_{1}|$, the presence of gradient staleness increases the escape speed from the saddle point. In contrast, when $|v_{2}|>|v_{1}|$, the equation $v_{1}^{2}(1+v_{1}\tau)^{2}+v_{2}^{2}(1+v_{2}\tau)^{2}=v_{1}^{2}+v_{2}^{2}$ has a single non-zero positive root, at which point the escape velocity from the saddle point matches that of the non-stale case.
Next, we have the following discussion on the characteristic root when $\tau\gg 1$.
Proposition 7
For $\tau\gg 1$, it holds that $\tilde{\lambda}=\frac{\log(-v\tau)-\log\log(-v\tau)}{\tau}+\mathcal{O}\left(\frac{\log\log(-v\tau)}{\tau\log{-v\tau}}\right)$.
Lastly, we consider the dynamics of ASGD along the eigenvector where the eigenvalue is zero, i.e., $v=0$. By Eq. (12), the dynamics of ASGD evolve according to
$\mathrm{d}x(t)=-q\mathrm{d}t+\sigma\mathrm{d}B(t),$ (34)
which is a Brownian motion with drift. In this case, it can be seen that the gradient staleness has no impact on the training process by Eq. (34). Next, without loss of generality, we focus on the case $q\leq 0$. In this case, for a Brownian motion starting at $x_{0}$, let us next present the first hitting time distribution of Eq. (34) to $x_{0}+\delta$, where $\delta>0$ is a constant. More specifically, the first hitting time is defined as $H_{0}=\inf\{t\geq 0:x(t)\geq x_{0}+\delta,x(0)=x_{0},q<0\}$, where $x(t)$ evolves according to Eq. (34). Note that $H_{0}$ is the first hitting time to $[x_{0}+\delta,\infty)$, or equivalently, the first exit time from $(-\infty,x_{0}+\delta)$. In this case, the probability distribution of $H_{0}$ is given by $p_{H_{0}}(t)=\frac{\delta}{\sqrt{2\pi t^{3}}}\exp{-\frac{(\delta-qt)^{2}}{2t}},~{}t>0.$
Next, the objective variable evolves according to
$\mathrm{d}x(t)=-v(1+v\tau)x(t)\mathrm{d}t+{(1+v\tau)}\sigma\mathrm{d}B(t),$ (35)
for $\tau\ll 1$ by [64]. In this way, we have the following remark for the connection between ASGD with stale gradients and SGD with momentum.
Remark:  From Eq. (28), we observe that a small degree of staleness can accelerate the convergence. Specifically, the drift coefficient is $-v(1+v\tau)$ in Eq. (35) with a gradient staleness $\tau$, while the drift coefficient is $-v$ in Eq. (17) without gradient staleness. This indicates that the introduction of gradient staleness increases the learning rate by a factor of $(1+\tau v)$. A careful reader may also observe a connection between small gradient staleness and the stochastic gradient descent method with momentum. Both approaches utilize information from previous gradients, which contributes to improved convergence rates. In particular, the formula of SGD with momentum is given by
$\breve{\boldsymbol{x}}[j+1]=\breve{\boldsymbol{x}}[j]-\eta\left(\nabla f(\breve{\boldsymbol{x}}[j])\right)+\beta(\breve{\boldsymbol{x}}[j]-\breve{\boldsymbol{x}}[j-1]).$ (36)
By [55, Section 7.3.1] and [65, Chapter 8.3.2], if the momentum algorithm always observes a gradient $\boldsymbol{g}$, then
$\Delta{\boldsymbol{x}}[j]=-\frac{\eta}{1-\beta}\boldsymbol{g}.$ (37)
where $\Delta{\boldsymbol{x}}[j]=\breve{\boldsymbol{x}}[j+1]-\breve{\boldsymbol{x}}[j]$. More specifically, the result of the momentum term is to increase the effective learning rate from $\eta$ to $\eta/(1-\beta)$ [55]. In this context, a small degree of gradient staleness and SGD with momentum both increase the convergence rate by leveraging the information of previous gradients. This effect is to increase the effective learning rate parameter. Further, we can express both gradient update methods in a unified form as follows:
$\Delta{\boldsymbol{x}}[j]=\sum_{l=0}^{j-1}\eta_{l}\left(\nabla f(\breve{\boldsymbol{x}}[j-1])\right)+\beta\Delta{\boldsymbol{x}}[j-l].$ (38)
For SGD with momentum, we set $\eta_{0}=\eta$ and $\beta_{1}=\beta$, while all other parameters are zero. For ASGD with a constant staleness $\breve{\tau}$, we have $\eta_{\breve{\tau}}=\eta$, with all other parameters remaining zero.
@@
In the following, we present the performance analysis when the gradient staleness is a uniform random variable, which can arise in event-triggered SGD. More specifically, consider the delay differential equation given below:
@@ IV Parameter Optimization for Distributed SGD
IV-A Parameter Optimization for ASGD
In this subsection, we present the probability distribution of the update interval, also referred to as the run-time[16], and step staleness in $B$-ASGD to bridge the ASGD algorithm and the SDDE-based continuous approximation. Recall that in $B$-ASGD, the central server collects the first $B$ gradients while the other clients do not halt the computation of their gradient. For the sake of discussion, we assume infinite communication bandwidth in this subsection, in which the communication delays are negligible compared to the computational delay. In this case, the gradient delay or the continuous-time gradient staleness equals the computational delay. The computational delay of each client is assumed to follow a general distribution, denoted by $T_{\kappa}$, where $\kappa$ represents the client index. Let $N_{p}(t)$ be the aggregated arrival process of gradients to the central server in $B$-ASGD. Let $Y\sim\text{Exp}(\nu)$ denote a random variable following an exponential distribution with a mean $1/\nu$, i.e., a rate $\nu$.
Let $N_{\kappa}(t)$ be the $\kappa$-th gradient arrival process, which is defined as the number of gradients collected by the central server until time $t$ from the $\kappa$-th worker. Clearly, $N_{\kappa}(t)$ is a step process, where the increments $N_{\kappa}(t)-N_{\kappa}(s)$, take on only non-negative integral values. In addition, a step process $N(t)$ is a Poisson process when $\text{Pr}\{N(t)-N(s)=k\}=\exp\{-(\Lambda(t)-\Lambda(s))\}\frac{(\Lambda(t)-\Lambda(s))^{k}}{k!},$ where $\Lambda(t)$ is the rate function of the Poisson process, $k\in\mathbb{N}_{0}$, and $\mathbb{N}_{0}$ is the set of non-negative integers. Let the aggregated process of $K$ step processes be denoted by $N_{p}^{K}(t)=\sum_{\kappa=1}^{K}N_{\kappa}(t)$. Here, $N_{p}^{K}(t)$ has the same meaning as $N_{p}(t)$ defined above. Here, we use the superscript $K$ when it is necessary to emphasize that $N_{p}^{K}(t)$ is a superposition of $K$ step processes. Next, let us introduce the following notations. Let $p_{\kappa}(k;t,s)=\Pr\{N_{\kappa}(t)-N_{\kappa}(s)\}=k$ for $s<t$ and $k\in\mathbb{N}_{0}$. Let $\Lambda_{K}(t,s)=\sum_{\kappa=1}^{K}p_{\kappa}(1;t,s)$ and $\Omega_{K}(t,s)=\sum_{\kappa=1}^{K}(1-p_{\kappa}(0;t,s)-p_{\kappa}(1;t,s))$. Next, we make the following assumption on the step processes [66, Theorem 30.3].
Assumption 3
It is assumed that the gradient arrival processes satisfy (1) $\lim_{K\to\infty}\Lambda_{K}(t,s)=\Lambda(t)-\Lambda(s)$, (2) $\lim_{K\to\infty}\Omega_{K}(t,0)=0$, and (3) $\lim_{K\to\infty}\max_{1\leq\kappa\leq K}(1-p_{\kappa}(0;t,0))=0$.
In the following, we provide a justification for Assumption 3, highlighting its practical relevance in the context of large-scale distributed learning. First, the Poisson approximation holds under the condition that the number of participating sub-processes tends to infinity, which aligns with the fact that a large number of workers are involved in the training process in large-scale distributed learning. Second, it is required that no two events (i.e., gradients from different clients) arrive simultaneously, which is reasonable as the probability of such simultaneous arrivals is practically zero. Third, for $N_{p}(t)$ with a rate $\Lambda_{K}$, the arrival rate of each sub-process should remain of lower order than $\mathcal{O}(\Lambda_{K})$ as $K\to\infty$, which implies that the gradient contribution from any single client does not dominate the overall process. Furthermore, in the case of a bandwidth-limited communication channel with a fixed service rate, to ensure the stability of the queues and prevent their divergence, the aggregate gradient arrival rate is a pre-given constant or function with respect to time, as presented in Assumption 3, and will be detailed in Subsection IV-D.
@@ IV-A Parameter Optimization for ASGD
Under Assumption 3, the aggregated gradient arrival process in pure ASGD $N_{p}^{K}(t)$ converges to a Poisson process with a rate function $\Lambda_{K}(t)$ as $K\to\infty$ in case $N_{\kappa}(t)$ are mutually independent.
Furthermore, [67, Theorem 2] presents an error analysis for approximating $N_{p}^{K}(t)$ by a Poisson process with finite $K$. This result demonstrates that, even if the individual step processes do not converge to a limiting process, the finite sum can nonetheless be well-approximated by a Poisson process.
In this way, we have the following corollary, which does not rely on exponentially distributed computation time [68, 69, 70].
Corollary 7
The update interval in $B$-ASGD is given by
@@
in which $I^{(i)}\sim\text{Exp}((K-i+1)\nu)$ and $\mathbb{E}\{I\}=\frac{1}{v}\sum_{i=K-B+1}^{K}\frac{1}{i}$ in case $\mathbb{E}\{T_{\kappa}\}=1/\nu$ for $\kappa\in\{1,2,...,K\}$. Also, it holds that
@@
By the Poisson approximation of the aggregated arrival process, the update interval between the arrival of the $(i-1)$th and $i$th gradients is exponentially distributed with rate $(K-i+1)\nu$ for $i\in\{1,...,B\}$. By this means, we get Eq. (42). Next, we have Eq. (43) by the Euler–Maclaurin formula, where the approximation error is on the order of $\mathcal{O}(1/{K^{2}})$. ∎
In synchronous SGD, the parameter server updates the global model only after receiving all the gradients from the workers. Therefore, the time between two consecutive updates in synchronous SGD is given by
$T_{s}^{\text{SYN}}=\max\{\hat{\tau}_{1},...,\hat{\tau}_{K}\},$ (44)
where $\hat{\tau}_{\kappa}$ represents the continuous-time gradient delay of Worker $\kappa$.
Remark:  To illustrate the straggler problem in synchronous SGD, we compare it with how asynchronous SGD addresses this issue. We focus on a scenario with $K=2$ workers as an example, each experiencing an exponentially computational delay. Additionally, the communication delays are assumed to be negligible. Note that in this case the continuous-time gradient delay is identical to the computational delay. Let $\hat{\tau}_{1}\sim\text{Exp}(\nu_{1})$ and $\hat{\tau}_{2}\sim\text{Exp}(\nu_{2})$, where it is assumed that $\nu_{1}\leq\nu_{2}$ without loss of generality. In this context, Work 1 is the straggler. The rate of the aggregated gradient arrival process in ASGD is $\nu_{p}=\nu_{1}+\nu_{2}$. By Eq. (44), we have $T_{s}^{\text{SYN}}=\max\{\hat{\tau}_{1},\hat{\tau}_{2}\}$. Let $\nu_{p}^{\text{SYN}}=\frac{1}{\mathbb{E}\{T_{s}^{\text{SYN}}\}}$. Therefore, $\nu_{p}^{\text{SYN}}=\frac{2}{\frac{1}{\nu_{1}}+\frac{1}{\nu_{2}}-\frac{1}{\nu_{1}+\nu_{2}}}$. It can be verified that $\nu_{p}>\nu_{p}^{\text{SYN}}$. The reason for this result is that the central server has to wait for the slowest worker before it can update the global model in synchronous distributed learning. Further, it holds that $\nu_{p}^{\text{SYN}}=\frac{2}{\frac{1}{v_{1}}+\frac{v_{1}}{v_{2}{(v_{1}+v_{2})}}}$. Hereby, we see that $\nu_{p}^{\text{SYN}}\approx 2\nu_{1}$, when $\nu_{2}\gg\nu_{1}$ as the contribution from the term $\frac{v_{1}}{v_{2}{(v_{1}+v_{2})}}$ becomes negligible compared to that of $\frac{1}{v_{1}}$. This illustrates that synchronous SGD is slowed down by the straggler worker.
@@
The expectation of the step staleness in $B$-ASGD without gradient dropout is given by [^10] When $B=1$, this theorem reduces to the pure ASGD case, which was presented by Eq. (18) in [17].
@@
It is observed that $B$-ASGD, the reduction of the step staleness is achieved by decreasing the update frequency. Next, we present a brief discussion on the difference between $B$-ASGD and $B$-batch ASGD. Similar to Theorem 7, the update interval follows an Erlang distribution, denoted by $\text{Erlang}(B,\nu)$, with mean $B/\nu$ in $B$-batch ASGD. Further, following the discussion in Theorem 2, it holds that $\mathbb{E}\{\breve{\tau}\}=K-1$ in $B$-batch ASGD.
To further reduce the step staleness, a client could choose to stop the computation of the current gradient according to a continuous-time or discrete-time staleness threshold. Next, we first present the following discussion on $B$-ASGD with a continuous-time staleness threshold. In the rest of this subsection, we assume each client’s computational delay follows an i.i.d. distribution, denoted by $T$. Let $F_{T}(\cdot)$ be the cumulative distribution function of $T$. In this context, let $p=F_{T}(\hat{\tau}_{th})$ be the probability that a computation of gradient is not discarded due to timeout. Further, $\bar{p}=1-p$. Let $\tilde{T}$ be the truncated computation time, in which $F_{\tilde{T}}(s)=\frac{F_{T}(s)}{p}$ if $s<\hat{\tau}_{th}$ and $F_{\tilde{T}}(s)=0$ for $s>\hat{\tau}_{th}$. In this case, $\mathbb{E}\{\tilde{T}\}=\frac{1}{p}\int_{s=0}^{\hat{\tau}_{th}}s{\rm d}{F_{T}(s)}$. Next, we have the following proposition on the inter-update time of the $B$-ASGD with a continuous-time staleness threshold.
@@
Next, we turn our attention to $B$-ASGD with a discrete-time gradient staleness threshold $\breve{\tau}_{th}$. In particular, if the global model has been updated $\breve{\tau}_{th}$ times before a client finishes its computation of the current gradient, the client shall stop the computation of the current gradient and start the computation of a new one. In this case, the step staleness of each gradient is no greater than $B\breve{\tau}_{th}$. With memoryless computation time, we have the following proposition.
@@
$\text{Pr}\{\breve{\tau}=mB\}=\frac{q\bar{q}^{m}}{1-\bar{q}^{\breve{\tau}_{th}+1}}\text{ for }\breve{\tau}=0,B,...,B\breve{\tau}_{th},$ (47)
@@
IV-B Parameter Optimization for Pure ASGD
In this subsection, we focus on pure ASGD, or $1$-ASGD, for deeper insights, in which $B=1$ and $\hat{\tau}_{th}=\infty$. Here we relax the assumption of identically distributed computation time across each client. Let $T_{\kappa}$ be the continuous-time random variable corresponding to the computational delay of the $\kappa$th worker. Further, it is assumed that $0<\mathbb{E}\{T_{\kappa}\}<\infty$. Let $\nu_{\kappa}=\frac{1}{\mathbb{E}\{T_{\kappa}\}}$ and $\nu_{p}=\sum_{\kappa=1}^{K}\nu_{\kappa}$. The sequence of the computational delay from Worker $\kappa$ is denoted by $\left(T_{\kappa}^{(i)}\right)_{i\geq 1}$. The update time of each worker can therefore be denoted by $J^{(n)}_{\kappa}=\sum_{i=1}^{n}T_{\kappa}^{(i)}$. Let us define the following stochastic process $N_{\kappa}(t)=\sum_{n=1}^{\infty}\mathds{1}\left\{J^{(n)}_{\kappa}\leq t\right\}$ representing the number of updates from Worker $\kappa$ by time $t$. Furthermore, the rate of $N_{\kappa}(t)$ is $\nu_{\kappa}$. In this context, the number of arrived gradients at the parameter server is given by the superposition of $N_{\kappa}(t)$ for $\kappa=1,2,...,K$. More specifically, we have $N_{p}(t)=\sum_{\kappa=1}^{K}N_{\kappa}(t).$ From Theorem 2, we directly have the following corollary.